# On a pinning model in correlated Gaussian random environments

### v1

@@ 3. The Skorohod case
In Section 3.1, we study the weak convergence for the Wick ordered partition function defined in (1.10) and prove Theorem 1.1 which guarantees the weak convergence of the rescaled Wick ordered partition functions in $L^{2}$-regime when $H\in(\frac{1}{2},1)$ and $\alpha>\frac{1}{2}$. In Section 3.2, we discuss the uniform integrability and characterize the weak limit of the Wick ordered partition functions for the case $\alpha\leq\frac{1}{2}$.
By Taylor expansion and Wick exponential (2.6), we get
@@ 3.1. Weak convergence
We shall focus on the case $11-H<{\alpha}<$ 1 and the proofs for the other two cases are similar.
@@
$\|\psi_{m,k}\|_{|\mathcal{H}|^{\otimes m}}^{2}\leq C_{H}^{m}\|\phi_{m+k}\|_{L^{1/H}(\mathbb{R}^{m+k})}^{2}\leq\frac{C^{m+k}[(m+k)!]^{2H}}{\Gamma\left((m+k)\frac{\alpha+H-1}{H}+1\right)^{2H}}.$
@@
and then it suffices to show $Y_{m,k}^{(N)}\rightarrow 0$ in $L^{2}$ as $N\rightarrow\infty$. By Lemma 2.4 and noting the symmetry of $f_{m,k}^{(N)}$ and $\psi_{m,k}$, we get
$\mathbb{E}\Big{[}\Big{(}Y_{m,k}^{(N)}\Big{)}^{2}\Big{]}\leq m!C^{m}\|f_{m,k}^{(N)}-\psi_{m,k}\|_{|\mathcal{H}|^{\otimes m}}^{2}.$
Thus, it remains to show
$\lim_{N\to\infty}\|f_{m,k}^{(N)}-\psi_{m,k}\|_{|\mathcal{H}|^{\otimes m}}^{2}=0.$
Define
@@
Together with (1.6) and (1.7), we have that ($t_{i}$ is unnecessarily belonging to $\Omega_{N}$).
@@
and thus by Lemma A.4, $\|f_{m,k}^{(N)}-\psi_{m,k}\|_{|\mathcal{H}|^{\otimes m}}^{2}\leq C_{H}^{m}\|f_{m,k}^{(N)}-\psi_{m,k}\|_{L^{1/H}(\mathbb{R}^{m})}^{2}\to 0$, as $N\to\infty$. ∎
@@
We assume (A1) and $\alpha>\frac{1}{2}$, i.e., $H\in(\frac{1}{2},1)$ and $\alpha>\frac{1}{2}$. Note that
@@
$=\sum_{k=0}^{\infty}\sum_{p+q=k}\frac{\hat{\beta}^{p}\hat{h}^{q}}{p!q!}\mathbf{I}_{p}\bigg{(}\int_{[0,1]^{q}}\phi_{k}(t_{1},\cdots,t_{k}){\mathrm{d}}t_{p+1}\cdots{\mathrm{d}}t_{k}\bigg{)}$
$=\sum_{k=0}^{\infty}\sum_{p+q=k}\frac{\hat{\beta}^{p}\hat{h}^{q}}{p!q!}\mathbf{I}_{p}(\psi_{p,q})=\sum_{m=0}^{\infty}\sum_{k=0}^{\infty}\frac{\hat{\beta}^{m}\hat{h}^{k}}{m!k!}\mathbf{I}_{m}(\psi_{m,k}),$
@@
By Proposition 3.1, we have that $\tilde{Z}_{N}^{(R)}\overset{(d)}{\longrightarrow}\tilde{Z}^{(R)}$ as $N\to\infty$. To complete the proof, we will apply Lemma A.3. Hence, it suffices to prove, as $R\rightarrow\infty$, $\tilde{Z}^{(R)}\overset{L^{2}}{\longrightarrow}\tilde{Z}_{\hat{\beta},\hat{h}}$ and $\tilde{Z}_{N}^{(R)}\overset{L^{2}}{\longrightarrow}\tilde{Z}_{N}$ uniformly in $N$, which is done in the following two steps respectively.
@@
$\Big{\|}\tilde{Z}^{(R)}-\tilde{Z}_{\hat{\beta},\hat{h}}\Big{\|}_{L^{2}}$ $=\Big{\|}\sum_{k+m>R}\frac{\hat{\beta}^{m}\hat{h}^{k}}{m!k!}\mathbf{I}_{m}^{(N)}(f_{m,k}^{(N)})\Big{\|}_{L^{2}}\leq\sum_{k+m>R}\frac{\hat{\beta}^{m}\hat{h}^{k}}{m!k!}\Big{\|}\mathbf{I}_{m}^{(N)}(f_{m,k}^{(N)})\Big{\|}_{L^{2}}$
$\leq C^{m+k}\sum_{k+m>R}\frac{\hat{\beta}^{m}\hat{h}^{k}}{\sqrt{m!}k!}\|\psi_{m,k}\|_{|\mathcal{H}|^{\otimes m}}\leq C_{H}^{m+k}\sum_{k+m>R}\frac{\hat{\beta}^{m}\hat{h}^{k}}{\sqrt{m!}k!}\|\phi_{m+k}\|_{L^{1/H}(\mathbb{R}^{m+k})}$
$\leq\sum_{k+m>R}(C_{\hat{\beta},\hat{h}}^{\alpha,H})^{m+k}\frac{[(m+k)!]^{H}}{\sqrt{m!}k!\Gamma\left((m+k)\frac{\alpha+H-1}{H}+1\right)^{H}},$
where the second inequality is due to the Riemann sum approximation and the third inequality follows from Lemma 3.2. By Stirling’s formula, each term in the sum on the right-hand side is bounded above by
$C^{m+k}\frac{(m+k)^{H/2}}{\sqrt{m^{1/2}k}\left((m+k)\frac{\alpha+H-1}{H}\right)^{H/2}}\frac{(m+k)^{(m+k)H}}{m^{m/2}k^{k}(m+k)^{(m+k)(\alpha+H-1)}}\leq C^{m+k}\frac{(m+k)^{(m+k)(1-\alpha)}}{m^{m/2}k^{k}}.$ (3.13)
@@
$C^{m+k}\Big{(}\frac{m^{(1-\alpha)m}m^{(1-\alpha)k}}{m^{m/2}k^{k}}+\frac{k^{(1-\alpha)k}k^{(1-\alpha)m}}{m^{m/2}k^{k}}\Big{)},$
@@
$\begin{split}&\Big{\|}\tilde{Z}^{(R)}_{N}-\tilde{Z}_{N}\Big{\|}_{L^{2}}^{2}=\mathbb{E}\Big{[}\Big{(}\sum_{k+m>R}\frac{1}{m!k!}S_{m,k}^{(N)}\Big{)}^{2}\Big{]}\\ =&\sum\limits_{k+m>R}\sum\limits_{k^{\prime}+m>R}\frac{\hat{h}^{k+k^{\prime}}\hat{\beta}^{2m}}{k!(k^{\prime})!(m!)^{2}}\mathbb{E}\left[\mathbf{I}_{m}^{(N)}(f_{m,k}^{(N)})\mathbf{I}_{m}^{(N)}(f_{m,k^{\prime}}^{(N)})\right]\\ \leq&\sum\limits_{k+m>R}\sum\limits_{k^{\prime}+m>R}\frac{\hat{h}^{k+k^{\prime}}\hat{\beta}^{2m}}{k!(k^{\prime})!(m!)^{2}}C^{m}m!\|f_{m,k}^{(N)}\|_{|\mathcal{H}|^{\otimes m}}\|f_{m,k^{\prime}}^{(N)}\|_{|\mathcal{H}|^{\otimes m}}.\end{split}$ (3.14)
@@
where the inequality follows from Cauchy-Schwartz inequality and Lemma 2.4. Recalling (3.9) and (3.10), by Jensen’s inequality and Lemma A.4, then (3.14) is bounded above by
@@
$\begin{split}&\sum\limits_{k+m>R}\sum\limits_{k^{\prime}+m>R}\frac{C^{2m+k+k^{\prime}}}{k!(k^{\prime})!m!}\|g_{m,k}^{(N)}\|_{L^{1/H}(\mathbb{R}^{m+k})}\|g_{m,k^{\prime}}^{(N)}\|_{L^{1/H}(\mathbb{R}^{m+k})}\\ \leq&\sum\limits_{k+m>R}\sum\limits_{k^{\prime}+m>R}\frac{C^{2m+k+k^{\prime}}}{k!(k^{\prime})!m!}\Big{\|}\phi_{m+k}^{\frac{1-{\alpha}+\epsilon}{1-{\alpha}}}\Big{\|}_{L^{1/H}(\mathbb{R}^{m+k})}\Big{\|}\phi_{m+k^{\prime}}^{\frac{1-{\alpha}+\epsilon}{1-{\alpha}}}\Big{\|}_{L^{1/H}(\mathbb{R}^{m+k})}\\ \leq&\sum_{k+m>R}C^{m+k}\frac{\hat{\beta}^{m}\hat{h}^{k}}{\sqrt{m!}k!}\Big{\|}\phi_{m+k}^{\frac{1-{\alpha}+\epsilon}{1-{\alpha}}}\Big{\|}_{L^{1/H}(\mathbb{R}^{m+k})}\sum_{k^{\prime}+m^{\prime}>R}C^{m^{\prime}+k^{\prime}}\frac{\hat{\beta}^{m^{\prime}}\hat{h}^{k^{\prime}}}{\sqrt{m^{\prime}!}k^{\prime}!}\Big{\|}\phi_{m^{\prime}+k^{\prime}}^{\frac{1-{\alpha}+\epsilon}{1-{\alpha}}}\Big{\|}_{L^{1/H}(\mathbb{R}^{m^{\prime}+k^{\prime}})},\end{split}$
@@
where the second inequality follows from (3.12) and the term on the right-hand side goes to 0 as $R\to\infty$ by choosing $\epsilon$ sufficiently small such that $(\alpha-\epsilon)+H>1$ and ${\alpha}-\epsilon>\frac{1}{2}$ and the result in Step (1). ∎
@@
Note that $\alpha=\frac{1}{2}$ is indeed a criticality. For simplicity, assume $h_{N}\equiv 0$ and $L(\cdot)\equiv 1$ in (1.11). Denote by $\tau^{\prime}$ an independent copy of $\tau$. Then
$\begin{split}\mathbb{E}\Big{[}\big{(}\tilde{Z}_{N}\big{)}^{2}\Big{]}&=\mathbb{E}_{\tau,\tau^{\prime}}\Big{[}\exp\Big{\{}\beta_{N}^{2}\sum\limits_{n,m=1}^{N}\gamma(n-m)\mathbf{1}_{\{n\in\tau\}}\mathbf{1}_{\{m\in\tau^{\prime}\}}\Big{\}}\Big{]}\\ &\geq\exp\Big{\{}\beta_{N}^{2}\sum\limits_{n,m=1}^{N}\gamma(n-m)\Big{\}}{\mathbb{P}}(\{1,\cdots,N\}\subset\tau\cap\tau^{\prime})\\ &\geq\exp\Big{(}C\hat{\beta}^{2}\frac{1}{N^{2(\alpha+H-1)}}\sum\limits_{n,m=1}^{N}|n-m|^{2H-2}\Big{)}({\mathbb{P}}(\tau_{1}=1))^{2N}\\ &\geq\exp(CN^{2(1-\alpha)})({\mathbb{P}}(\tau_{1}=1))^{2N},\end{split}$
which tends to infinity as $N\to\infty$ if $\alpha<\frac{1}{2}$. In this case, $\{\tilde{Z}_{N}\}$ is not $L^{2}$ uniformly bounded.

### v2

@@ 3. The Skorohod case
In Section 3.1, we study the weak convergence for the Wick-ordered partition function defined in (1.11) and prove Theorem 1.1 which guarantees the weak convergence of the rescaled Wick-ordered partition functions in the $L^{2}$-regime when $H\in(\frac{1}{2},1)$ and $\alpha\geq\frac{1}{2}$. In Section 3.2, we discuss the uniform integrability and characterize the weak limit of the Wick-ordered partition functions for the case $\alpha\leq\frac{1}{2}$.
By a Taylor expansion and Wick exponential (2.8), we get
@@ 3.1. Weak convergence
Moreover, for any $p\in\mathbb{N}$ and $m_{1},\cdots,m_{p},k_{1},\cdots,k_{p}\in\mathbb{N}_{0}$, we have the following joint convergence in distribution
$\left(\mathbf{I}_{m_{1}}^{(N)}(f_{m_{1},k_{1}}^{(N)}),\cdots,\mathbf{I}_{m_{p}}^{(N)}(f_{m_{p},k_{p}}^{(N)})\right)\overset{(d)}{\longrightarrow}\left(\mathbf{I}_{m_{1}}(\psi_{m_{1},k_{1}}),\cdots,\mathbf{I}_{m_{p}}(\psi_{m_{p},k_{p}})\right),\quad\text{as}~N\rightarrow\infty.$ (3.8)
@@
We shall focus on the case $1-H<{\alpha}<1$ and the proofs for the other two cases are similar.
@@
$\|\psi_{m,k}\|_{|\mathcal{H}|^{\otimes m}}^{2}\leq C_{H}^{m}\|\phi_{m+k}\|_{L^{1/H}(\mathbb{R}^{m+k})}^{2}\leq\frac{C^{m+k}[(m+k)!]^{2H}}{\Gamma\big((m+k)\frac{\alpha+H-1}{H}+1\big)^{2H}}.$
@@
and then it suffices to show $Y_{m,k}^{(N)}\rightarrow 0$ in $L^{2}$ as $N\rightarrow\infty$. By Lemma 2.5 and noting the symmetry of $f_{m,k}^{(N)}$ and $\psi_{m,k}$, we get $\mathbb{E}[(Y_{m,k}^{(N)})^{2}]\leq m!C^{m}\|f_{m,k}^{(N)}-\psi_{m,k}\|_{|\mathcal{H}|^{\otimes m}}^{2}.$ Thus, it remains to show $\lim_{N\to\infty}\|f_{m,k}^{(N)}-\psi_{m,k}\|_{|\mathcal{H}|^{\otimes m}}^{2}=0.$ Define
@@
Together with (1.6) and (1.7), we have that for $t_{i}\in(\frac{n_{i}-1}{N},\frac{n_{i}}{N}],i=1,\dots,m+k$,
@@
and thus by Lemma A.4, $\|f_{m,k}^{(N)}-\psi_{m,k}\|_{|\mathcal{H}|^{\otimes m}}^{2}\leq C_{H}^{m}\|f_{m,k}^{(N)}-\psi_{m,k}\|_{L^{1/H}(\mathbb{R}^{m})}^{2}\to 0$, as $N\to\infty$.
The joint convergence in distribution (3.8) is then a direct consequence of (2.17) in Proposition 2.7 and $Y_{m_{i},k_{i}}^{(N)}\to 0$ in $L^{2}$ as $N\to\infty$ above for each $i=1,\cdots,p$. ∎
@@
Case (1): We assume (A1) and $\alpha>\frac{1}{2}$, i.e., $H\in(\frac{1}{2},1)$ and $\alpha>\frac{1}{2}$. Note that
@@
[…]
@@
By (3.8) in Proposition 3.1 and the Cramér-Wold Theorem, we have that $\tilde{Z}_{N}^{(R)}\overset{(d)}{\longrightarrow}\tilde{Z}^{(R)}$ as $N\to\infty$. To complete the proof, we will apply Lemma A.3. Hence, it suffices to prove, as $R\rightarrow\infty$, $\tilde{Z}^{(R)}\overset{L^{2}}{\longrightarrow}\tilde{Z}_{\hat{\beta},\hat{h}}$ and $\tilde{Z}_{N}^{(R)}\overset{L^{2}}{\longrightarrow}\tilde{Z}_{N}$ uniformly in $N$, which is done in the following two steps respectively.
@@
$\Big\|\tilde{Z}^{(R)}-\tilde{Z}_{\hat{\beta},\hat{h}}\Big\|_{L^{2}}$ $=\Big\|\sum_{k+m>R}\frac{\hat{\beta}^{m}\hat{h}^{k}}{m!k!}\mathbf{I}_{m}(\psi_{m,k})\Big\|_{L^{2}}\leq\sum_{k+m>R}\frac{\hat{\beta}^{m}|\hat{h}|^{k}}{m!k!}\Big\|\mathbf{I}_{m}(\psi_{m,k})\Big\|_{L^{2}}$ (3.14)
$\leq C_{H}^{m+k}\sum_{k+m>R}\frac{\hat{\beta}^{m}|\hat{h}|^{k}}{\sqrt{m!}k!}\|\phi_{m+k}\|_{L^{1/H}(\mathbb{R}^{m+k})}$
$\leq\sum_{k+m>R}(C_{\hat{\beta},\hat{h}}^{\alpha,H})^{m+k}\frac{[(m+k)!]^{H}}{\sqrt{m!}k!\Gamma\big((m+k)\frac{\alpha+H-1}{H}+1\big)^{H}},$
where the second and the last inequality is due to Lemma 3.2. By Stirling’s formula, each term in the sum on the right-hand side is bounded above by
$C^{m+k}\frac{(m+k)^{H/2}}{\sqrt{m^{1/2}k}\left((m+k)\frac{\alpha+H-1}{H}\right)^{H/2}}\frac{(m+k)^{(m+k)H}}{m^{m/2}k^{k}(m+k)^{(m+k)(\alpha+H-1)}}\leq C^{m+k}\frac{(m+k)^{(m+k)(1-\alpha)}}{m^{m/2}k^{k}},$ (3.15)
@@
$C^{m+k}\Big(\frac{m^{(1-\alpha)m}m^{(1-\alpha)k}}{m^{m/2}k^{k}}+\frac{k^{(1-\alpha)k}k^{(1-\alpha)m}}{m^{m/2}k^{k}}\Big)$
@@
$\begin{split}\Big\|\tilde{Z}^{(R)}_{N}-\tilde{Z}_{N}\Big\|_{L^{2}}^{2}&=\mathbb{E}\Big[\Big(\sum_{k+m>R}\frac{1}{m!k!}S_{m,k}^{(N)}\Big)^{2}\Big]\\ &=\sum\limits_{k+m>R}\sum\limits_{k^{\prime}+m>R}\frac{\hat{h}^{k+k^{\prime}}\hat{\beta}^{2m}}{k!(k^{\prime})!(m!)^{2}}\mathbb{E}\left[\mathbf{I}_{m}^{(N)}(f_{m,k}^{(N)})\mathbf{I}_{m}^{(N)}(f_{m,k^{\prime}}^{(N)})\right]\\ &\leq\sum\limits_{k+m>R}\sum\limits_{k^{\prime}+m>R}\frac{|\hat{h}|^{k+k^{\prime}}\hat{\beta}^{2m}}{k!(k^{\prime})!(m!)^{2}}C^{m}m!\|f_{m,k}^{(N)}\|_{|\mathcal{H}|^{\otimes m}}\|f_{m,k^{\prime}}^{(N)}\|_{|\mathcal{H}|^{\otimes m}}.\end{split}$ (3.16)
@@
where the inequality follows from the Cauchy- Schwarz inequality and Lemma 2.5. Recalling (3.10) and (3.11), by Jensen’s inequality and Lemma A.4, then (3.16) is bounded above by
@@
$\begin{split}&\sum\limits_{k+m>R}\sum\limits_{k^{\prime}+m>R}\frac{C^{2m+k+k^{\prime}}}{k!(k^{\prime})!m!}\|g_{m,k}^{(N)}\|_{L^{1/H}(\mathbb{R}^{m+k})}\|g_{m,k^{\prime}}^{(N)}\|_{L^{1/H}(\mathbb{R}^{m+k})}\\ \leq&\sum\limits_{k+m>R}\sum\limits_{k^{\prime}+m>R}\frac{C^{2m+k+k^{\prime}}}{k!(k^{\prime})!m!}\Big\|\phi_{m+k}^{\frac{1-{\alpha}+\epsilon}{1-{\alpha}}}\Big\|_{L^{1/H}(\mathbb{R}^{m+k})}\Big\|\phi_{m+k^{\prime}}^{\frac{1-{\alpha}+\epsilon}{1-{\alpha}}}\Big\|_{L^{1/H}(\mathbb{R}^{m+k})}\\ \leq&\sum_{k+m>R}\frac{C^{m+k}}{\sqrt{m!}k!}\Big\|\phi_{m+k}^{\frac{1-{\alpha}+\epsilon}{1-{\alpha}}}\Big\|_{L^{1/H}(\mathbb{R}^{m+k})}\sum_{k^{\prime}+m^{\prime}>R}\frac{C^{m^{\prime}+k^{\prime}}}{\sqrt{m^{\prime}!}k^{\prime}!}\Big\|\phi_{m^{\prime}+k^{\prime}}^{\frac{1-{\alpha}+\epsilon}{1-{\alpha}}}\Big\|_{L^{1/H}(\mathbb{R}^{m^{\prime}+k^{\prime}})}.\end{split}$
@@
where the first inequality follows from (3.13). By the same estimate as in (3.14) above, the right-hand side of the last inequality goes to 0 as $R\to\infty$ by choosing $\epsilon$ sufficiently small such that $(\alpha-\epsilon)+H>1$ and ${\alpha}-\epsilon>\frac{1}{2}$.
Case (2): We assume (A1) and $\alpha=\frac{1}{2}$, i.e., $H\in(\frac{1}{2},1)$ and $\alpha=\frac{1}{2}$. The proof is almost the same as in Case (1) and we indicate only the necessary modifications. First, noting that by the assumption $0<\inf_{n\in\mathbb{N}}L(n)\leq\sup_{n\in\mathbb{N}}L(n)<+\infty$ for this case, the exponent $\epsilon$ in (3.12) vanishes. Then by (3.14), (3.15) and (3.16), it suffices to show that
$\sum\limits_{m,k=0}^{\infty}\big(\hat{\beta}C_{H}\big)^{m}\big(|\hat{h}|C_{H}\big)^{k}\frac{[(m+k)!]^{H}}{\sqrt{m!}k!\Gamma((m+k)(\frac{2H-1}{2H})+1)^{H}}\leq\sum\limits_{m,k=0}^{\infty}\big(\hat{\beta}C_{H}\big)^{m}\big(|\hat{h}|C_{H}\big)^{k}\frac{(m+k)^{\frac{m+k}{2}}}{m^{m/2}k^{k}}$ (3.17)
converges for $\hat{\beta}>0,|\hat{h}|>0$ small enough.
Note that
$\frac{(m+k)^{\frac{m+k}{2}}}{m^{m/2}k^{k}}\leq\frac{(2k)^{k}}{m^{m/2}k^{k}}\mathrm{I}_{\{k\geq m\}}+\frac{(2m)^{\frac{m+k}{2}}}{m^{m/2}k^{k}}\mathrm{I}_{\{k\leq m\}}\leq\frac{2^{k}}{m^{m/2}}\mathrm{I}_{\{k\geq m\}}+\frac{2^{m}m^{\frac{k}{2}}}{k^{k}}\mathrm{I}_{\{k\leq m\}}.$
Then the right-hand side of (3.17) is bounded by
$\begin{split}&\sum\limits_{m=0}^{\infty}\frac{(\hat{\beta}C_{H})^{m}}{m^{m/2}}\sum\limits_{k=m}^{\infty}(2|\hat{h}|C_{H})^{k}+\sum\limits_{m=0}^{\infty}(2\hat{\beta}C_{H})^{m}\sum\limits_{k=0}^{m}\frac{(C_{H}|\hat{h}|\sqrt{m})^{k}}{k!}\\ \leq&\sum\limits_{m=0}^{\infty}\frac{(\hat{\beta}C_{H})^{m}}{m^{m/2}}\cdot\frac{1}{1-2|\hat{h}|C_{H}}+\sum\limits_{m=0}^{\infty}e^{C_{H}|\hat{h}|\sqrt{m}}(2C_{H}\hat{\beta})^{m}<+\infty,\end{split}$
given that $2C_{H}|\hat{h}|<1$ and $2C_{H}\hat{\beta}<1$.
Case (3): We assume $\alpha<\frac{1}{2}$. Denote by $\tau^{\prime}$ an independent copy of $\tau$ and we have that
$\begin{split}\mathbb{E}\Big[\tilde{Z}_{N}^{2}\Big]&=\mathbb{E}_{\tau,\tau^{\prime}}\Big[\exp\Big\{\beta_{N}^{2}\sum\limits_{n,m=1}^{N}\gamma(n-m)\mathbf{1}_{\{n\in\tau\}}\mathbf{1}_{\{m\in\tau^{\prime}\}}+h_{N}\sum\limits_{n=1}^{N}\big(\mathbf{1}_{\{n\in\tau\}}+\mathbf{1}_{\{n\in\tau^{\prime}\}}\big)\Big\}\Big]\\ &\geq\exp\Big\{\beta_{N}^{2}\sum\limits_{n,m=1}^{N}\gamma(n-m)-2N|h_{N}|\Big\}{\mathbb{P}}(\{1,\cdots,N\}\subset\tau\cap\tau^{\prime})\\ &\geq\exp\Big(C\hat{\beta}^{2}\frac{1}{N^{2(\alpha+H-1+\epsilon)}}\sum\limits_{n,m=1}^{N}|n-m|^{2H-2}-C|\hat{h}|N^{1-\alpha+\epsilon}\Big)({\mathbb{P}}(\tau_{1}=1))^{2N}\\ &\geq\exp\Big(C\Big(N^{2(1-\alpha-\epsilon)}-N^{1-\alpha+\epsilon}\Big)\Big)({\mathbb{P}}(\tau_{1}=1))^{2N},\end{split}$
which tends to infinity as $N\to\infty$ if $\alpha<\frac{1}{2}$, by choosing $\epsilon>0$ small enough in (3.12). ∎
@@
We now comment on the case $\alpha=\tfrac{1}{2}$, which represents the threshold between the $L^{2}$ regime and the non-$L^{2}$ regime. However, this should not be regarded as the marginal case between the disorder-relevant and disorder-irrelevant regimes, since in Section 3.2 we show that the subsequential limit of $\{\tilde{Z}_{N}\}_{N}$ can be nonzero even for some $\alpha<\tfrac{1}{2}$. Moreover, the case $\alpha=\tfrac{1}{2}$ can still be treated within the framework of [1, 9], where disorder is relevant, rather than that of [10], where disorder is marginally relevant. These two frameworks are fundamentally different.
In comparison with the parabolic Anderson model on $\mathbb{R}^{2}$, where the $L^{2}$ solution exists only locally while the $L^{1}$ solution exists globally (see [28]), we believe that when $\alpha=\tfrac{1}{2}$, the $L^{2}$ regime of our model exists only locally and the $L^{1}$ regime exists globally (see Remarks 1.41.6, and Section 3.2) We also note that our assumption on $L(\cdot)$ may not be optimal. To allow for a general slowly varying function $L(\cdot)$, one would need to verify the dominated convergence theorem carefully. Since our main purpose here is to highlight the critical nature of $\alpha=\tfrac{1}{2}$ by a simple argument, we are content with the present restriction on $L(\cdot)$.